A 20 kg block is in motion on a rough horizontal surface.a horizontal force of 60 N is required to keep the body moving with constant speed.the coefficient of kinetic friction is

Grade:9

2 Answers

Arun
5 years ago

m=20 kg

F=μmg

or μs=F/mg

=60/20*98

=60/196=0.306

 

 

Hope it helps

 

 

Regards

Arun (askIITians forum expert)

Kartik
5 years ago
 

m=20 kg

F=μmg

or μs=F/mg

=60/20*98

=60/196=0.306

Hence u can do all questions like these 

Regards and wish u the best for jee

Thnks

 

Hope it helps